OpenStudy (anonymous):

4x-4y=10 3x+2y=5 my options: (2,-1/2) (1,1) (3,-2)

hero (hero):

Multiply the second equation by 2, then add both equations together: 4x - 4y = 10 6x + 4y = 10 10x = 20

hero (hero):

Finish solving from there

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

(2,1/2)?

OpenStudy (anonymous):

*(2,-1/2)

hero (hero):

Yeah, and substitute x = 2 back into the original equation to find y.

OpenStudy (anonymous):

I got it(: thanks!

OpenStudy (nurali):

Let 4x - 4y = 10........(i) 3x + 2y = 5.........(ii) equation (i) divide by '2' and Add (i) and (ii) 2x - 2y = 5 3x + 2y = 5 ----------- 5x = 10 x = 10/5 = 2 Substituting the value of x in (ii) 3 * 2 + 2y = 5 2y = 5 - 6 y = - 1/2 Therefore, x = 2 and y = - 1/2